# Optimizing Continuous-Wave Pumped Entanglement-based QKD in Noisy Environment

### v2

@@ II.2 Coincidence Measurement in Noisy Settings
| Dark count | 25 cps |
II.3 Characterization of Detector Parameters
The SPCM-AQRH-10 detector is a commonly used single-photon detection module in QKD. This is an actively quenched Geiger mode detector, and the manufacturer-provided parameters of this detector are summarized in Table. 1. To study the count rate-dependent detector parameters, from the experimental setup described in Fig. 1, we only considered the APD2 output and investigated detector parameters such as detector dead time and twilight events. For this, the singles values recorded from APD2 were subtracted from their neighboring photon time values, that is, where is the time of arrival of photons and is the index value. The number of measured time separations () is plotted as a histogram in Fig. 3. This figure gives interesting results about many detector parameters, including dead-time and twilight events. The dead time in single-photon avalanche photo detectors [21] is often described as a temporal window in which photons cannot be detected, which means that the dead time can be estimated by knowing the separation from 0 to the offset value of time in Fig. 3.
We observed that the dead time is not constant but increases with the count rate. This observation contradicts the common assumption that detector dead time is constant. An increasing dead time with higher count rates suggests that the detector recovery dynamics is more complex. Under high illumination, the internal circuitry may take more time to fully reset, possibly due to saturation effects or an increased thermal load.
The varying detector dead time can influence the temporal shift of the coincidence histogram discussed in the Section II.2. That is, the dead time can be defined as an interval that spans the quenching and resetting time. In this interval, there is a time zone called the twilight zone, where the detector is semi-sensitive, and photons may be sensed but will not be reflected until the detector completes resetting. This means that the photons in the twilight zone must wait for a period of time to be registered. This waiting period is responsible for a delay in the cross-correlation function. The dead time and twilight events are shown in Fig. 3 in two cases of noise values. As illustrated, the twilight events occur close to the dead time of the detectors because if two photons arrive at the detectors closer to the dead time, the lagging one will have to wait for the dead time to end or will not be registered at all. If a large number of photons are sent to the detectors, this scenario will happen more frequently, and so the twilight event becomes significantly high. The measured values of the twilight events and the dead time for various noise levels are shown in Fig. 3 in panels (c) and (d). The change in twilight events and dead time increased sharply after the noise level passed 4.5 Mcps. This is in proportion to an observed shift in the coincidence histogram where shift and FWHM also drastically changed after reaching this onset value.
@@ III Theoretical Analysis
Given the trend of coincidence events in the presence of extreme noise, one should investigate its influence on the SKR and QBER. Typically, a coincidence window (a chosen measurement time window to observe coincidence events) is optimized to achieve a maximum SKR with a manageable QBER. Due to the Gaussian shape of coincidence events, optimal coincidence window selection is crucial for improved QKD performance. Additionally, the change in the shape of the coincidence histogram due to extreme noise should also be considered when choosing the coincidence window.
@@ III.1 Coincidence Measurement in a Single Basis System
Eq. (9) provides a complete measurement result, including the noise effect on the coincidence profile. Therefore, the total coincidence values within the chosen coincidence window can be estimated by carrying out the integration
Where is the total number of coincidence events with start-bin and stop-bin . When we determine the start-bin and stop-bin values and compute the total coincidence events, , using Eq. (10), the result will contain all coincidence events including coincidence events from generated photon pairs and accidental coincidence events. Also, it will have a detector efficiency factor, as our model has accounted for it. When the detectors operating in a low noise regime, the share of accidental coincidence in the computed “C” value is low. In this case, it is desirable to select a wide coincidence window. However, in extreme noise events, the share of accidental coincidence is higher, and the target in this case would be to minimize the accidental coincidences so that QBER can be lowered. The QBER and accidental coincidence are directly proportional and therefore, in a single basis system, minimizing the ratio of accidental coincidence with total coincidence would be one target parameter that could approximate QBER measurement, i.e.
Fig. 4 shows the values of for different start and stop bin values, where the coincidence window size is defined as the difference between the stop and start bins (stop bin - start bin). When noise is low, reaches desirable low levels, with an observed value of = 0.1% in this case. As noise increases, the values of increase accordingly to 7. 59%, 15. 32%, and 19. 9% for Figs. 4 (b), (c), and (d), respectively.
Since the width of the coincidence histogram increases with noise, different start and stop bin values must be considered to achieve a lower . We selected the following start and stop bin values (in ns): 3 and 4.6, 3.75 and 5.05, 5 and 6.3, 5 and 6.3, respectively. Notably, the start bin value must be adjusted to correct for the shift in the coincidence histogram’s peak position.
@@ III.2 QKD Performance Under Noisy Conditions
where ‘’ and ‘’ indicate the positions of the detectors in the experiment as depicted in Fig. 5. Eq. (13) will provide complete coincidence values generated with the flexibility to select the coincidence window. Notice that the value in Eq. (9) will turn to as total generated photon pairs must be equally shared between detectors. With this, the QBER () of the system can be defined as
@@
When implementing the protocol, Alice and Bob randomly switch their measurement basis between and . Next, they publicly reveal their basis choice. Due to random selection, they have a 50% probability of choosing the correct basis. Only compatible basis choices are considered for protocol implementation, hence, 50% of the coincidence measurements are discarded. Therefore, the total achievable secure key rate () in this procedure could be given by [30]
@@
Using Eq. (18) and (14), we examine SKR and QBER in noisy scenarios. With a pair rate () of 500K cps distributed equally at the Alice and Bob sites, Fig. 6 illustrates the trend of SKR and QBER in various noise levels for different start-bin and stop-bin choices. In this demonstration, we considered equal noise distributed to all detectors. The result shows that a QBER threshold of (indicated by the dark dotted line in Fig. 6) is established, above which no secure key can be generated. At this threshold, the secure key rate becomes zero (). Therefore, achieving is a target for the implementation of QKD. To reach this threshold, different start and stop bin choices tolerate different levels of noise. For example, when the start bin = 3 ns and the stop bin = 6 ns, the system tolerates 26 million noise counts before reaching . However, it accepts 17 million noise counts when the start of the bin = 1 ns and the stop bin = 8 nm, despite the larger total coincidence window. When the start bin is set near the centroid of the coincidence window (the start bin = 4.5 ns) and the stop bin = 7.5 ns, the system shows reduced resistance to noise, tolerating only 13.5 million noise counts.
III.3 Optimization of Coincidence Window
To improve the secure key rate () value even in higher noise counts, we narrow the coincidence window by adjusting the start and stop bin (because of the Gaussian-like profile of the coincidence histogram). Taking this into account, we selected the start-bin () = 4 ns and the stop-bin () = 5 ns. As expected, the noise immunity for this choice showed better performance, requiring 32 Mcps to reach = 0. However, the problem with this choice of coincidence window is the reduced secure key rate when the noise is lower. To optimize the coincidence window, we apply the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm with the objective condition:
The objective function in our case is . Our aim is to maximize with fixed noise values, and the photon pair rate (). The negative value of indicates the maximization in BFGS, where and are the variables. Therefore, in the BFGS formulation, we take and . With the initialization value and the initial Hessian function approximation , we calculate the gradient in each iteration as and the search direction . The updates of the values of in each iteration can be found using , where is the step size. We compute the differences and and achieve the updated Hessian approximation as:
where . This process continues until reaches a value less than the error tolerance. With the given values of and , the BFGS algorithm can obtain a maximum value of . Our requirement of achieving the best possible value when the system is operating in a noisy environment can be satisfied. By assigning equal noise values to our model (with given total noise level Mcps) and setting Kcps, we found that with ns and ns, the system will give the best performance for the QKD system. With this choice of coincidence window, the system can be used to obtain 33,162 secure keys, reaching when Mcps.
Following this theoretical model, one could examine QKD in a variety of experimental settings, including cases of QKD execution under high pair rate and low noise, high pair rate and high noise, and low pair rate and high noise. Additionally, our model is suitable for the characterization of QKD systems in imbalanced detections where one detector receives many more photons than the others. This is especially the case in the entanglement-based QKD implementation where the source is on Alice’s side [15].
The case presented in this work considered the pair rate of 500 Kcps. With previous techniques [24], it is possible to achieve pair rates as high as 10 Mcps and could also maximize the pair rate by following special techniques using bulk crystals [32]. The optimum values of start and stop bins will be different for different pair rates. This demonstrates the importance of characterizing the system parameters before employing them in practical use.
@@ IV Conclusion
Furthermore, our demonstrated noise-induced shift in detection timing events possibly threatens the overall security of QKD, as it is well known that shifts in coincidence detection event times can be exploited by eavesdroppers for the hacking of secret data [33]. More studies are needed to understand the extent to which this shift influences the security of the system. One possible solution to mitigate the rate-induced timing shift is to characterize detector efficiency versus input photon rate analysis, as shown in [34], and then discard all communication events when the photon rate exceeds the onset value of efficiency degradation. From our analysis, efficiency degradation and coincidence histogram shift dominated after 4.5 Mcps (although the rated maximum count rate of the detector used is much higher, 17 Mcps).
@@
V Acknowledgment
The experiment was carried out by Hashir Kuniyil while working at Ozyegin University under a project funded by the Turkish Scientific and Technological Research Council (TÜBTAK) with Project No. 118E991. The idea development and all subsequent processing, including theoretical analysis and manuscript writing, were performed by the listed authors, including Hashir Kuniyil, at their respective institutions.

### v4

@@ II.3 Detection timing shift and FWHM widening in start-stop histogram
II.4 Characterization of Detector Parameters
The SPCM-AQRH-14 detector is a commonly used single-photon detection module in QKD [22, 32, 33]. This is an actively quenched Geiger mode detector and a nonparalyzable detector; after a detection event, it remains insensitive for a fixed dead time . The manufacturer-provided parameters of this detector are summarized in Table. 1. To demonstrate that the effective recovery time of the avalanche photodiode (APD) increases with photon count rate, we first characterized the timing response of a single detector operated under varying detection rates. From the time-tagged photon arrival records, we evaluated the temporal separation between successive detection events, defined as
Histograms of were constructed using a timing resolution of ps, as shown in Fig. 3. At low detection rates, the distribution of is approximately flat over the chosen time window, indicating a uniform probability of detecting photons at different separations (though minimum detection time is found out to be 22 ns, exactly the dead time value provided by the manufacturer). This behavior reflects the fact that the detector has sufficient time to fully recover between successive detection events. As the photon count rate increases, the probability of observing short inter-arrival times increases noticeably, with a pronounced accumulation of events near the minimum detectable separation, as shown in Fig. 3(b).
We interpret this minimum photon separation as the effective detector dead time, defined as the shortest interval following a detection during which the detector can register a subsequent photon. With further increases in the photon count rate, detection events at this minimum separation occur with increasing frequency, providing direct evidence that the effective dead time—or, more generally, the detector recovery dynamics—becomes rate dependent as the detector is driven toward saturation.
The extracted effective dead time as a function of the photon count rate is summarized in Fig. 4. To describe this behavior, the data were fitted using a saturating recovery model of the form
where is the initial (or baseline) delay time, is the maximum additional delay induced by recovery effects at high count rates, and is a characteristic recovery rate. The fit captures the observed increase and saturation of the effective dead time with increasing photon flux.
To establish that these recovery dynamics are responsible for the shift and broadening observed in the coincidence histograms, we compare the extracted single-detector recovery behavior with two-detector coincidence measurements. As shown in Fig. 2(b), the coincidence peak position shifts systematically with increasing noise count rate, while Fig. Fig. 2(b) also shows a concomitant increase in the coincidence peak full width at half maximum (FWHM). Both quantities are well described by the same functional dependence on the count rate, with the coincidence peak shift following Eq. 15. and the FWHM following an analogous saturating form. Importantly, the characteristic rate extracted independently from the single-detector dead time analysis (Fig. 3), the coincidence peak shift, and the coincidence peak broadening (Fig. 2(b)) is consistent across all measurements. This quantitative agreement demonstrates that the observed coincidence peak shift and broadening originate from the same underlying physical mechanism: rate-dependent detector recovery dynamics in the APD. The consistency between single-detector timing statistics and two-detector coincidence measurements confirms that detector recovery effects, rather than optical or electronic artifacts, govern the observed distortions in the coincidence profiles.
@@ III Theoretical Analysis
Given the trend of coincidence counts in the presence of extreme noise, one should investigate its influence on the SKR and QBER. Typically, a coincidence window (a chosen measurement time window to observe coincidence counts) is optimized to achieve a maximum SKR with a manageable QBER. Due to the Gaussian shape of coincidence counts, optimal coincidence window selection is crucial for improved QKD performance. Additionally, the change in the shape of the coincidence histogram due to extreme noise should also be considered when choosing the coincidence window.
@@ III.1 Coincidence Measurement in a Single Basis System
Eq. (9) provides a complete measurement result, including the noise effect on the coincidence profile. As a result, the total coincidence values within the chosen coincidence window can be estimated by carrying out the integration
Where is the total number of coincidence counts with start-bin and stop-bin . When we determine the start-bin and stop-bin values and compute the total coincidence counts, , using Eq. (20), the result will contain all coincidence counts including coincidence counts from generated photon pairs and accidental coincidence counts. Notice that our experiment is carried out with time tagger resolution of 85 ps. Because the experimental histogram is reported in counts per bin per second, we evaluate the total coincidence rate in a window by summing over bins (equivalently, integrating the corresponding rate density over) in Eq. (20). When the detectors operate in a low noise regime, the share of accidental coincidence in the computed “C” value is low. In this case, it is desirable to select a wide coincidence window. However, in extreme noise events, the share of accidental coincidence is higher, and the target in this case would be to minimize the accidental coincidences so that QBER can be lowered. The QBER and accidental coincidence are directly proportional and hence, in a single basis system, minimizing the ratio of accidental coincidence with total coincidence would be one target parameter(accidental coincidence fraction, ) that could approximate QBER measurement, i.e.
where is the estimate of accidental coincidence, experimentally we can estimate it by , where accidental coincidence per bin, () is the number of bins selected, and is the experimental duration. From our measured and fitted curve we can estimate :
In QKD, this is closely related to an error contribution, but whether it equals the QBER depends on the encoding/basis. If accidentals are uncorrelated with the correct outcome, they typically contribute error among the accidental subset. which means .
Since the width of the coincidence histogram increases with noise plus the its timing shift essentially require us to optimize the coincidence window. It is clear from our experimental results that fixed choice of coincidence window will cause loss of useful measurement and its wrong choice would lead to noise induced high QBER.
@@ III.2 QKD Performance Under Noisy Conditions
where ‘’ and ‘’ indicate the positions of the detectors in the experiment as depicted in Fig. 5. Eq. (24) will provide complete coincidence values generated with the flexibility to select the coincidence window. Notice that the brightness ,, in Eq. (9) will turn to as total generated photon pairs must be equally shared between detectors. With this, the QBER () of the system can be redefined as
@@
When implementing the protocol, Alice and Bob randomly switch their measurement basis between and . Next, they publicly reveal their basis choices. Due to random selection, they have only 50% probability of choosing the correct basis. Only compatible basis choices are considered for protocol implementation, hence, 50% of the coincidence measurements are discarded. Therefore, the total achievable secure key rate () in this procedure could be given by [34]
@@
Using Eq. (31) and (26), we examine SKR and QBER in noisy scenarios. With a pair rate () of 500K cps distributed equally at the Alice and Bob sites, Fig. 6 illustrates the trend of SKR and QBER in various noise levels for different start-bin and stop-bin choices. In this demonstration, we considered equal noise distributed to all detectors. The result shows that a QBER threshold of (indicated by the dark dotted line in Fig. 6) is established, above which no secure key can be generated. At this threshold, the secure key rate becomes zero (). Therefore, achieving is a target for the implementation of QKD. To reach this threshold, different start and stop bin choices tolerate different levels of noise. For example, when the start bin = 3 ns and the stop bin = 6 ns, the system tolerates 26 Mcps noise counts before reaching . However, it accepts 17 Mcps noise counts when the start of the bin = 1 ns and the stop bin = 8 ns, despite the larger total coincidence window. When the start bin is set near the centroid of the coincidence window (the start bin = 4.5 ns) and the stop bin = 7.5 ns, the system shows reduced resistance to noise, tolerating only 13.5 Mcps.
III.3 Optimization of the Coincidence Window
The coincidence window plays a critical role in determining the secure key rate (SKR) and quantum bit error rate (QBER), particularly in noisy operating conditions. While a narrower coincidence window reduces accidental coincidences and improves QBER at high noise levels, it also discards genuine coincidence events and degrades SKR in low-noise conditions. Therefore, an optimal choice of the coincidence window must balance these competing effects.
To identify the optimal coincidence window for a given photon pair rate and noise level, we numerically optimized the secure key rate with respect to the start and stop times of the coincidence window, using the theoretical model developed in Sec. III. The optimization was performed by maximizing under fixed experimental parameters, including the photon pair rate and detector noise levels.
Specifically, we employed a gradient-based optimization algorithm (Broyden–Fletcher–Goldfarb–Shanno, BFGS) to determine the coincidence window that maximizes the SKR for a given noise condition. The optimization procedure treats the coincidence window boundaries as free variables while all detector and source parameters are held fixed.
As a representative example, for a photon pair rate of kcps and a total noise level of Mcps distributed equally among the detectors, the optimization yields an optimal coincidence window of ns and ns. With this choice, the system achieves a maximum secure key rate of 33,162 bits and tolerates noise levels up to Mcps before the SKR drops to zero. The corresponding trends in SKR and QBER for different coincidence window choices are shown in Fig. 6.
@@ IV Conclusion
Furthermore, our demonstrated noise-induced shift in detection timing events possibly threatens the overall security of QKD, as it is well known that shifts in coincidence detection event times can be exploited by eavesdroppers for the hacking of secret data [36]. More studies are needed to understand the extent to which this shift influences the security of the system. One possible solution to mitigate the rate-induced timing shift is to characterize detector efficiency versus input photon rate analysis, as shown in [37], and then discard all communication events when the photon rate exceeds the onset value of efficiency degradation. From our analysis, we observed the onset of efficiency degradation and coincidence histogram shift after Mcps (although the rated maximum count rate of the detector used is much higher, Mcps). is the value manufacturer tested the dead time. From our method we found that default dead time of the detector is ns. this is the exact dead time provided by the manufacturer.
@@
Appendix A Delay-Scan Coincidence Histogram from Time-Tagged Events
Coincidence measurements were computed from raw time-tagged detection events using a delay-scan algorithm. The input data consist of a sequence of time tags and detector identifiers stored in a CSV file with the structure:
Time, Detector
where Time denotes the recorded time tag (integer-valued) and Detector specifies the detector/channel index.
The coincidence curve is obtained by applying a discrete delay to all timestamps from detector 1 and counting the number of events that coincide with timestamps from detector 2. Formally,
where and denote the sets of timestamps recorded by detector 1 and detector 2, respectively, and is the indicator function.
Important note. This implementation defines coincidences using exact timestamp equality. This approach is valid when timestamps are discretized into fixed-width time bins and coincidence is defined by bin alignment. For experiments requiring a finite coincidence window (e.g., ), the matching condition should be modified accordingly.
The following Python script was used to compute the coincidence histogram:
This procedure yields the coincidence distribution as a function of relative detector delay, allowing identification of the temporal correlation peak.
Appendix B Numerical Optimization Procedure
The optimization of the coincidence window boundaries was carried out using the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm, a quasi-Newton method for unconstrained optimization. The objective function was defined as the negative of the secure key rate, , such that minimizing corresponds to maximizing the SKR.
The optimization variables were the start and stop times of the coincidence window. Gradients were evaluated numerically, and the Hessian matrix was updated iteratively following the standard BFGS update rule. The optimization was terminated when the gradient norm fell below a predefined tolerance. This procedure was repeated for different noise levels and photon pair rates.